the parks commission wants a new rectangular sign with an area of 45 square meters for the visitor center. the length of the sign is to be 4 meters longer than the width. what will be the length and width of the new sign?
Answer by checkley77(12844)

L=W+4
W(W+4)=45
W^2+4W-45=0
(W+9)(W-5)=0
W-5=0
W=5 ANSWER.
L=5+4=9 FOR THE LENGTH.
PROOF:
5*9=45
45=45